\begin{thebibliography}{10}

\bibitem{Abu-Sini}
Maria Abu-Sini and Eitan Yaakobi.
\newblock On levenshtein’s reconstruction problem under insertions,
  deletions, and substitutions.
\newblock {\em IEEE Transactions on Information Theory}, 67(11):7132--7158,
  2021.

\bibitem{Bornholt}
James Bornholt, Randolph Lopez, Douglas~M. Carmean, Luis Ceze, Georg Seelig,
  and Karin Strauss.
\newblock A dna-based archival storage system.
\newblock In {\em Proceedings of the Twenty-First International Conference on
  Architectural Support for Programming Languages and Operating Systems},
  ASPLOS '16, page 637–649, New York, NY, USA, 2016. Association for
  Computing Machinery.

\bibitem{Gabrys}
Ryan Gabrys and Eitan Yaakobi.
\newblock Sequence reconstruction over the deletion channel.
\newblock {\em IEEE Transactions on Information Theory}, 64(4):2924--2931,
  2018.

\bibitem{Horovitz}
Michal Horovitz and Eitan Yaakobi.
\newblock Reconstruction of sequences over non-identical channels.
\newblock {\em IEEE Transactions on Information Theory}, 65(2):1267--1286,
  2019.

\bibitem{LEVENSHTEIN2001310}
Vladimir I.~Levenshtein.
\newblock Efficient reconstruction of sequences from their subsequences or
  supersequences.
\newblock {\em Journal of Combinatorial Theory, Series A}, 93(2):310--332,
  2001.

\bibitem{Levenshtein}
Vladimir I.Levenshtein.
\newblock Efficient reconstruction of sequences.
\newblock {\em IEEE Transactions on Information Theory}, 47(1):2--22, 2001.

\bibitem{Junnila}
Ville Junnila, Tero Laihonen, and Tuomo Lehtilä.
\newblock On levenshtein’s channel and list size in information retrieval.
\newblock {\em IEEE Transactions on Information Theory}, 67(6):3322--3341,
  2021.

\bibitem{Junnila2}
Ville Junnila, Tero Laihonen, and Tuomo Lehtilä.
\newblock The levenshtein’s sequence reconstruction problem and the length of
  the list.
\newblock {\em IEEE Transactions on Information Theory}, 70(2):1050--1066,
  2024.

\bibitem{JUNNILA2025115279}
Ville Junnila, Tero Laihonen, and Tuomo Lehtilä.
\newblock Levenshtein's sequence reconstruction problem and results for larger
  alphabet sizes.
\newblock {\em Theoretical Computer Science}, 1045:115279, 2025.

\bibitem{Liron}
Yuvalal Liron and Michael Langberg.
\newblock A characterization of the number of subsequences obtained via the
  deletion channel.
\newblock {\em IEEE Transactions on Information Theory}, 61(5):2300--2312,
  2015.

\bibitem{PHAM2025105980}
Van Long~Phuoc Pham, Keshav Goyal, and Han~Mao Kiah.
\newblock Sequence reconstruction problem for deletion channels: A complete
  asymptotic solution.
\newblock {\em Journal of Combinatorial Theory, Series A}, 211:105980, 2025.

\bibitem{Sala}
Frederic Sala, Ryan Gabrys, Clayton Schoeny, and Lara Dolecek.
\newblock Exact reconstruction from insertions in synchronization codes.
\newblock {\em IEEE Transactions on Information Theory}, 63(4):2428--2445,
  2017.

\bibitem{Hirschberg}
Daniel S.Hirschberg and Mireille Regnier.
\newblock Tight bounds on the number of string subsequences.
\newblock {\em Journal of Discrete Algorithms}, 1(1):123--132, 2000.

\bibitem{Yaakobi}
Eitan Yaakobi and Jehoshua Bruck.
\newblock On the uncertainty of information retrieval in associative memories.
\newblock {\em IEEE Transactions on Information Theory}, 65(4):2155--2165,
  2019.

\bibitem{Yaakoni2}
Eitan Yaakobi, Jehoshua Bruck, and Paul~H. Siegel.
\newblock Constructions and decoding of cyclic codes over $b$ -symbol read
  channels.
\newblock {\em IEEE Transactions on Information Theory}, 62(4):1541--1551,
  2016.

\bibitem{Yazdi}
S.~M. Hossein~Tabatabaei Yazdi, Han~Mao Kiah, Eva Garcia-Ruiz, Jian Ma, Huimin
  Zhao, and Olgica Milenkovic.
\newblock Dna-based storage: Trends and methods.
\newblock {\em IEEE Transactions on Molecular, Biological, and Multi-Scale
  Communications}, 1(3):230--248, 2015.

\end{thebibliography}

\documentclass[reqno,11pt]{amsart}

\usepackage{amssymb,amsmath,amsthm,amsfonts,mathtools,bbm,mathrsfs,enumitem,graphics,float,multirow}
\usepackage[normalem]{ulem}
\usepackage{hyperref}
\usepackage{comment}
\usepackage{hyperref}
\hypersetup{
    colorlinks,
    linkcolor={blue!80!black},
    citecolor={red!80!black},
    urlcolor={red!80!black}
}
\usepackage{xargs}                      
\usepackage{xcolor}  
\usepackage[colorinlistoftodos,prependcaption,textsize=tiny]{todonotes}
\newcommandx{\unsure}[2][1=]{\todo[linecolor=red,backgroundcolor=red!25,bordercolor=red,#1]{#2}}
\newcommandx{\change}[2][1=]{\todo[linecolor=blue,backgroundcolor=orange!25,bordercolor=blue,#1]{#2}}
\newcommandx{\info}[2][1=]{\todo[linecolor=OliveGreen,backgroundcolor=OliveGreen!25,bordercolor=OliveGreen,#1]{#2}}

\usepackage[top=1.6in,bottom=1.5in,left=1.3in,right=1.3in]{geometry}
\allowdisplaybreaks
\usepackage{cellspace}
\newcommand\redout{\bgroup\markoverwith{\textcolor{red}{\rule[0.5ex]{2pt}{0.8pt}}}\ULon}

\newtheorem{theorem}{Theorem}
\newtheorem{lemma}[theorem]{Lemma}

\usepackage{tikz}
\usetikzlibrary{calc}

\tikzset{font= }

\newcommand\nc\newcommand
\nc\bfa{{\boldsymbol a}}\nc\bfA{{\boldsymbol A}}\nc\cA{{\mathscr A}}
\nc\bfb{{\boldsymbol b}}\nc\bfB{{\boldsymbol B}}\nc\cB{{\mathscr B}}
\nc\bfc{{\boldsymbol c}}\nc\bfC{{\boldsymbol C}}\nc\cC{{\mathscr C}}
\nc\bfd{{\boldsymbol d}}\nc\bfD{{\boldsymbol D}}\nc\cD{{\mathscr D}}
\nc\bfe{{\boldsymbol e}}\nc\bfE{{\boldsymbol E}}\nc\cE{{\mathscr E}}
\nc\bff{{\boldsymbol f}}\nc\bfF{{\boldsymbol F}}\nc\cF{{\mathscr F}}
\nc\bfg{{\boldsymbol g}}\nc\bfG{{\boldsymbol G}}\nc\cG{{\mathscr G}}
\nc\bfh{{\boldsymbol h}}\nc\bfH{{\boldsymbol H}}\nc\cH{{\mathscr H}}\nc\fH{{\mathfrak H}}
\nc\bfi{{\boldsymbol i}}\nc\bfI{{\boldsymbol I}}\nc\cI{{\mathcal I}}
\nc\bfj{{\boldsymbol j}}\nc\bfJ{{\boldsymbol J}}\nc\cJ{{\mathscr J}}
\nc\bfk{{\boldsymbol k}}\nc\bfK{{\boldsymbol K}}\nc\cK{{\mathscr K}}
\nc\bfl{{\boldsymbol l}}\nc\bfL{{\boldsymbol L}}\nc\cL{{\mathscr L}}
\nc\bfm{{\boldsymbol m}}\nc\bfM{{\boldsymbol M}}\nc\cM{{\mathscr M}}
\nc\bfn{{\boldsymbol n}}\nc\bfN{{\boldsymbol N}}\nc\sN{{\mathscr N}}
\nc\bfo{{\boldsymbol o}}\nc\bfO{{\boldsymbol O}}\nc\cO{{\mathscr O}}
\nc\bfp{{\boldsymbol p}}\nc\bfP{{\boldsymbol P}}\nc\cP{{\mathscr P}}
\nc\bfq{{\boldsymbol q}}\nc\bfQ{{\boldsymbol Q}}\nc\cQ{{\mathscr Q}}
\nc\bfr{{\boldsymbol r}}\nc\bfR{{\boldsymbol R}}\nc\cR{{\mathscr R}}
\nc\bfs{{\boldsymbol s}}\nc\bfS{{\boldsymbol S}}\nc\cS{{\mathscr S}}
\nc\bft{{\boldsymbol t}}\nc\bfT{{\boldsymbol T}}\nc\cT{{\mathscr T}}
\nc\bfu{{\boldsymbol u}}\nc\bfU{{\boldsymbol U}}\nc\cU{{\mathscr U}}
\nc\bfv{{\boldsymbol v}}\nc\bfV{{\boldsymbol V}}\nc\cV{{\mathscr V}}
\nc\bfw{{\boldsymbol w}}\nc\bfW{{\boldsymbol W}}\nc\cW{{\mathscr W}}
\nc\bfx{{\boldsymbol x}}\nc\bfX{{\boldsymbol X}}\nc\cX{{\mathscr X}}
\nc\bfy{{\boldsymbol y}}\nc\bfY{{\boldsymbol Y}}\nc\cY{{\mathscr Y}}
\nc\bfz{{\boldsymbol z}}\nc\bfZ{{\boldsymbol Z}}\nc\cZ{{\mathscr Z}}
\nc\pp{\mathbb{P}}
\nc\ee{\mathbb{E} }

\renewcommand{\le}{\leqslant}
\renewcommand{\leq}{\leqslant}
\renewcommand{\ge}{\geqslant}
\renewcommand{\geq}{\geqslant}

\nc{\Cay}{{\sf Cay}}

\nc{\ff}{{\mathbb F}}

\newcommand\remove[1]{}

\title[]{Sequence Reconstruction over the Deletion Channel}
\author[]{Fengxing Zhu}\thanks{Institute for Systems Research and Department of ECE, University of Maryland, College Park, MD 20742, USA, fengxing@terpmail.umd.edu. Supported in part by NSF grant CCF 2330909.}
\date{}

\begin{document}
\begin{abstract}
In this paper, we consider the Levenshtein's sequence reconstruction problem in the case where the transmitted codeword is chosen from $\{0,1\}^n$ and the channel can delete up to $t$ symbols from the transmitted codeword. We determine the minimum number of channel outputs (assuming that they are distinct) required to reconstruct a list of size $\ell-1$ of candidate sequences, one of which corresponds to the original transmitted sequence. More specifically,  we determine the maximum possible size of the intersection of 
$\ell \geq 3$ deletion balls of radius $t$ centered at $x_1, x_2, \dots, x_{\ell}$, where 
$x_i \in \{0,1\}^n$ for all $i \in \{1,2,\dots,\ell\}$ and $x_i \neq x_j$ for 
$i \neq j$, with $n \geq t+ \ell-1$ and $t \geq 1$.
\end{abstract}
\maketitle

\section{Introduction}
The study of sequence reconstruction was initiated by Levenshtein in 
\cite{Levenshtein,LEVENSHTEIN2001310}, where a sender transmits a codeword $x$ over multiple noisy channels. The receiver observes the outputs of these channels and attempts to uniquely reconstruct the transmitted codeword $x$. For the fixed codebook $\mathcal{C}$ and the channel the main task is to determine the minimum number of channel outputs required to guarantee unique reconstruction. The motivation for the sequence reconstruction problem originates from biology and chemistry, where traditional redundancy-based error correction methods are unsuitable. In recent years, the problem has regained attention due to its strong relevance to information retrieval in advanced storage technologies. In such systems, the stored data may consist of a single copy that is read multiple times or several redundant copies of the same information~\cite{Horovitz,Yaakoni2}. This problem~\cite{Horovitz} is particularly significant in the context of DNA data storage~\cite{Yazdi,Bornholt}, where numerous noisy copies of DNA strands are available, and the objective is to accurately reconstruct the original information from these imperfect observations.

 This sequence reconstruction problem has been extensively studied under various channels. Hirschberg and Regnier in \cite{Hirschberg} derived tight bounds on the number of string subsequences. In \cite{Levenshtein,LEVENSHTEIN2001310}, Levenshtein obtained the minimum number of channel outputs for the deletion and insertion channel required for unique reconstruction for the case where $\mathcal{C}$ consists of all binary vectors of length $n$. Gabrys and Yaakobi~\cite{Gabrys} later solved the sequence reconstruction problem over the $t$-deletion channel, where $\mathcal{C}$ consists of binary vectors such that $d_L(x,y) \geq 2$ for $x,y \in \mathcal{C}$, with $d_L(x,y)$ being the Levenshtein distance between $x$ and $y$. More recently, Pham, Goyal, and Kiah~\cite{PHAM2025105980} obtained a complete asymptotic solution for this problem where $\mathcal{C}$ consists of binary vectors such that $d_L(x,y) \geq \ell$ for $x,y \in \mathcal{C}$.

The sequence reconstruction problem has also been studied under other channel models 
besides the deletion channel. Sala, Gabrys, Schoeny, and Dolecek~\cite{Sala} solved 
this problem for the insertion channel, assuming the codebook $\mathcal{C}$ consists 
of $q$-ary vectors satisfying $d_L(x,y) \geq \ell$ for all distinct $x, y \in \mathcal{C}$. 
Abu-Sini and Yaakobi~\cite{Abu-Sini} investigated the reconstruction problem 
for channels involving a single deletion combined with multiple substitutions, as well 
as channels involving a single insertion combined with a single substitution.

A variant of the sequence reconstruction problem allows the decoder to output a list 
of possible sequences instead of a unique reconstruction. Yaakobi and 
Bruck~\cite{Yaakobi} studied this problem for channels introducing substitution errors. 
In particular, they investigated the maximum intersection of $m$ Hamming balls of 
radius $t$ centered at $x_1, x_2, \dots, x_m$, where $d_H(x_i, x_j) \geq d$ for 
$i \neq j$. Junnila, Laihonen, and Lehtil\"a~\cite{Junnila,Junnila2} analyzed the 
list size when the channel introduces substitution errors with $t = e + \ell$, where 
$e$ is the error-correcting capability of a binary code $\mathcal{C}$. More recently, 
they extended these results from the binary case to the $q$-ary case 
in~\cite{JUNNILA2025115279}.

In this paper we focus on the deletion channel. Formally, when a a codeword of length $n$ is sent through a $t$-deletion channel, a subsequence of length $n-t$ is received. A $t$-deletion correcting code $\mathcal{C}$ is a subset of length-$n$ binary vectors such that for any vector $x \in \mathcal{C}$, $x$ can be uniquely identified from any length-$(n-t)$ subsequence of $x$. More specifically, we study the minimum number of $t$-deletion channel outputs (assuming they are distinct) required to reconstruct a list of size $\ell-1$ of candidate sequences, one of which corresponds to the original transmitted sequence when $\mathcal{C}=\{0,1\}^n$. In other words, we determine the maximum possible size of the intersection of 
$\ell \geq 3$ deletion balls of radius $t$ centered at $x_1, x_2, \dots, x_{\ell}$, where 
$x_i \in \{0,1\}^n$ for all $i \in \{1,2,\dots,\ell\}$ and $x_i \neq x_j$ for 
$i \neq j$, with $n \geq t+ \ell-1$ and $t \geq 1$.

\section{Definitions and Preliminaries}
Let $x$ be a binary sequence of length $n$ over $\mathbb{F}_2^n$. The deletion ball of radius $t$ centered at $x \in \mathbb{F}_2^n$ is define to be 
$$ \mathcal{D}_t(x)=\{y \in \mathbb{F}_2^{n-t} | y \textrm{ is a subsequence of } x \}.$$

For any two sequences $x_1$ and $x_2$, their Levenshtein distance is $t$ if $\mathcal{D}_t(x_1) \cap \mathcal{D}_t(x_2) \neq \emptyset$ and $\mathcal{D}_{t-1}(x_1) \cap \mathcal{D}_{t-1}(x_2) = \emptyset$

Let $a_n \in \mathbb{F}_2^n$ be an alternating sequences where its first bit is 1. For $0 < t < n$, we denote the maximum size of a deletion ball of radius $t$, by $D(n,t)$,i.e 
$$D(n,t)=\max_{x \in \mathbb{F}_2^n}|\mathcal{D}_t(x)|.$$
From \cite{Levenshtein} and \cite{LEVENSHTEIN2001310}, we know that 
$$D(n,t)=| \mathcal{D}_t(a_n)|=\sum_{i=0}^t \binom{n-t}{i}$$
and also 
$$D(n,t)=D(n-1,t)+D(n-2,t-1).$$
Note that $D(n,n)=1$ and $D(n,t)=0$ if $t <0$ or $n <t$. 

Due to Levenshtein in \cite{Levenshtein} and \cite{LEVENSHTEIN2001310}, we have 
$$\max_{x_1 \neq x_2 ,x_1,x_2 \in \mathbb{F}_2^n}|\mathcal{D}_t(x_1) \cap \mathcal{D}_t(x_2)|=2D(n-2,t-1).$$
In this paper we will study 
$$N(n,\ell,t):=\max_{x_1 \neq x_2 ...\neq x_{\ell} ; x_1,x_2,...,x_{\ell} \in \mathbb{F}_2^n}|\mathcal{D}_t(x_1) \cap \mathcal{D}_t(x_2) ...\cap \mathcal{D}_t(x_{\ell})|,$$
where $\ell \geq 3$, $n \geq t+ \ell-1$ and $t \geq 1$. Specifically, we establish the following theorem. 

\begin{theorem} \label{theorem: main}
For $\ell \geq 3$, $n \geq t+ \ell-1$ and $t \geq 1$, we have that 
\begin{align*}
    N(n,\ell,t)= \sum_{i=1}^{\ell-2}D(n-2i,t-i)+2D(n-2(\ell-1),t-(\ell-1)).
\end{align*}
\end{theorem}
We will adopt the techniques and analysis from \cite{Gabrys} and \cite{PHAM2025105980} to prove Theorem~\ref{theorem: main}. In particular, in order to prove the upper bound in Theorem~\ref{theorem: main}, we will use induction similar to that in \cite{Gabrys} and \cite{PHAM2025105980}. Specifically, we will prove it by induction on $n,\ell,t$. The case where $\ell=3$, $t \geq 1$ and $n \geq t+\ell-1$ will serve as part of the base case. We could use the case $\ell=2$ as a base case but in the proof of the general case there are certain places where $ \ell \geq 4$ is necessary and also we feel it is instructive to give the proof for $\ell=3$. 

Before proceeding, we need to give some definitions and state some lemmas that will be used very often in our analysis.

Let $\chi \subset \mathbb{F}_2^n$ be a set and $v$ a sequence of length at most $n$. We denote by $\chi^{v}$ the set of all sequences in $\chi$ that start with the sequence $v$, that is, 
$$\chi^{v}=\{x \in \chi | v \textrm{ is  a prefix of }x\}.$$

For a sequence $v \in \mathbb{F}_2^m$ and a set $\chi \in \mathbb{F}_2^n$, the set $v \circ \chi$ is prepending the sequence $v$ before every sequence in $\chi$, that is 
$$v \circ \chi=\{(vx) | x \in \chi \}.$$
The following two lemmas will be used very often in our analysis. Lemma \ref{lemma 1} was derived in \cite{Gabrys} and Lemma \ref{lemma 2} was obtained in \cite{Liron}. 

\begin{lemma} \label{lemma 1}
Let $n$, $m_1$, and $t$ be positive integers, and $x=x^1x^2\cdots x^n \in \mathbb{F}_2^n$, $x_1 \in \mathbb{F}_2^{m_1}.$ Assume that $k$ is the smallest integer such that $x_1$ is a subsequence of $(x^1,x^2,...,x^{k})$. Then 
$$\mathcal{D}_t(x)^{x_1}=x_1 \circ \mathcal{D}_{t^*}(x^{k+1},...,x^n),$$
where $t^*=t-(k-m_1).$ In particular,
$$|\mathcal{D}_t(x)^{x_1}|=|\mathcal{D}_{t^*}(x^{k+1},...,x^n)|.$$
\end{lemma}

 \begin{lemma} \label{lemma 2}
 Let $l <n$, $x \in \mathbb{F}_2^n$, $y \in \mathbb{F}_2^{n-l}$, where $y \in \mathcal{D}_{l}(x)$ and $l <t$. Then $\mathcal{D}_{t-l}(y) \subset \mathcal{D}_{t}(x)$.
\end{lemma}

For the ease of notation, let us define
$$N_{\ell}(n,t):=\sum_{i=1}^{\ell-2}D(n-2i,t-i)+2D(n-2(\ell-1),t-(\ell-1)).$$

\section{The Intersection of Three Deletion Balls}
In this section we will prove that 
$$N(n,3,t)=N_3(n,t).$$ 

\subsection{The Lower Bound}

In this section we will show that $3D(n-4,t-2)+D(n-3,t-1)$ is a lower bound for $N(n,3,t)$ by showing that sequences $x_1=10a_{n-2}$, $x_2=01a_{n-2}$ and $x_3=0101a_{n-4}$ satisfy 
$$|\mathcal{D}_t(x_1) \cap \mathcal{D}_t(x_2) \cap \mathcal{D}_t(x_3)|=3D(n-4,t-2)+D(n-3,t-1).$$
We have the following Theorem.
\begin{theorem}
For $t \geq 1$ and $n \geq t+2$,
$$N(n,3,t) \geq 3D(n-4,t-2)+D(n-3,t-1).$$
\end{theorem}

\begin{proof} 
Let $\chi=\mathcal{D}_t(x_1) \cap \mathcal{D}_t(x_2) \cap \mathcal{D}_t(x_3)$. If $a=a^1a^2\cdots a^j$ is binary sequence of length $j$, then we denote $\bar a=\bar a^1\bar a^2 \cdots \bar a^j$ as a sequence of length $j$ such that $\bar a^i= (1-a^i)$ for $i \in \{1,2\cdots,j\}$. By Lemma~\ref{lemma 1} and Lemma~\ref{lemma 2}, we have 
\begin{align*}
|\chi^{00}|& =|00 \circ \mathcal{D}_{t-2}(a_{n-4}) \cap \mathcal{D}_{t-2}(a_{n-4}) \cap \mathcal{D}_{t-1}(1a_{n-4})|\\
&=|\mathcal{D}_{t-2}(a_{n-4})|\\
&=D(n-4,t-2).
\end{align*}
\begin{align*} 
|\chi^{01}|& =|01 \circ \mathcal{D}_{t-1}(\bar a_{n-3}) \cap \mathcal{D}_{t}(a_{n-2}) \cap \mathcal{D}_{t}(01a_{n-4})|\\
&=|\mathcal{D}_{t-1}(\bar a_{n-3})|\\
&=D(n-3,t-1).
\end{align*}
\begin{align*}
|\chi^{11}|& =|11 \circ \mathcal{D}_{t-1}(\bar a_{n-3}) \cap \mathcal{D}_{t-1}(\bar a_{n-3}) \cap \mathcal{D}_{t-2}(a_{n-4})|\\
&=|\mathcal{D}_{t-2}(a_{n-4})|\\
&=D(n-4,t-2).
\end{align*}
\begin{align*}
|\chi^{10}|& =|10 \circ \mathcal{D}_{t}(a_{n-2}) \cap \mathcal{D}_{t-2}(a_{n-4}) \cap \mathcal{D}_{t-1}(1a_{n-4})|\\
&=|\mathcal{D}_{t-2}(a_{n-4})|\\
&=D(n-4,t-2).
\end{align*}

Since $\chi=\chi^{00} \cup \chi^{01} \cup \chi^{10} \cup \chi^{11}$, the proof is done. 
\end{proof}
\subsection{The Upper Bound}
 We will show that $N_3(n,t)$ is an upper bound for $N(n,3,t)$.

 We will prove it by induction on $n$ and $t$. Let us first address the base case. 

The base case is when $n=t+2$ and $t \geq 1$. since $D(n-2,t-1)=2$ and $D(n-4,t-2)=1$, $N_3(n,t)=4$. It is easy to see that after deleting $t$ symbols from each sequence of length $t+2$ there are only $2$ symbols left for each sequence. Since $N_3(n,t)=4$, $N(n,3,t) \leq N_3(n,t)$. 

Now let $t=1$ and $n \geq t+2$. Then we have $N_3(n,t)=1$. By \cite{LEVENSHTEIN2001310}, we know that every two distinct binary sequences of length $n$ can have at most $2$ common supersequences of length $2$. Therefore, we have $N(n,3,1) \leq 1=N_3(n,1)$. 

Now we will move onto the induction step. Assume that $N(n_0,3,t_0) \leq N_3(n_0,t_0)$ is true for all $n_0 \geq t_0+2$ and  $t_0\geq 1$ such that $n_0+t_0 < n+t$. We will need a few lemmas to complete this step.
 
\begin{lemma} \label{lemma: x_1^1,x_2^2,x_3^1}
Assume that $t \geq 1$ and $n \geq t+2$. Let $x_1$,$x_2$ and $x_3$ be three arbitrary sequences in $\mathbb{F}_2^n$ such that $x_1 \neq x_2 \neq x_3$ and $a=x_1^1=x_2^1=x_3^1$. Then 
$$|\mathcal{D}_t(x_1) \cap \mathcal{D}_t(x_2) \cap \mathcal{D}_t(x_3)|\leq N_3(n,t).$$
\end{lemma}

\begin{proof}Our proof follows the proof of Theorem 8 in \cite{Gabrys}. Let $\chi=\mathcal{D}_t(x_1) \cap \mathcal{D}_t(x_2) \cap \mathcal{D}_t(x_3).$ We have 
\begin{align*}
|\chi^{a}|& =|a \circ \mathcal{D}_{t}(x_1^2,...,x_1^n)\cap \mathcal{D}_{t}(x_2^2,...,x_2^n) \cap \mathcal{D}_{t}(x_3^2,...,x_3^n|\\
&\leq \max_{x \neq y \neq z,x,y,z \in \mathbb{F}_2^{n-1}}|\mathcal{D}_{t}(x)\cap \mathcal{D}_{t}(y) \cap \mathcal{D}_{t}(z)|\\
& \leq N_3(n-1,t).
\end{align*}

Suppose $x_1^k$ is the first occurrence of the symbol $\bar{a}$ in $x_1$ and the symbol $\bar{a}$ appears in $x_1$ not after it appears in $x_2$ and $x_3.$ If $x_1^k=x_2^k=x_3^k=\bar{a}$ we have 
\begin{align*}
|\chi^{\bar{a}}|& =|\bar{a} \circ \mathcal{D}_{t-(k-1)}(x_1^{k+1},...x_1^n)\cap \mathcal{D}_{t-(k-1)}(x_2^{k+1},...,x_2^n) \cap \mathcal{D}_{t-(k-1)}(x_3^{k+1},...,x_3^n)|\\
&\leq \max_{x \neq y \neq z,x,y,z \in \mathbb{F}_2^{n-k}}|\mathcal{D}_{t-(k-1)}(x)\cap \mathcal{D}_{t-(k-1)}(y) \cap \mathcal{D}_{t-(k-1)}(z)|\\
& \leq N_3(n-k,t-k+1)\\
& \leq N_3(n-2,t-1).
\end{align*}
If one of $x_2^k$ or $x_3^k$ is equal to $a$, say $x_2^k=a$ then we have 
\begin{align*}
|\chi^{\bar{a}}|& \leq |\bar{a} \circ  \mathcal{D}_{t-k}(x_2^{k+2},...x_2^n)|\\
&\leq D(n-k-1,t-k)\\
& \leq D(n-3,t-2)\\
& \leq N_3(n-2,t-1).
\end{align*}
Since $|\chi^a|+|\chi^{\bar{a}}|\leq N_3(n-1,t)+N_3(n-2,t-1)=N_3(n,t)$, the proof is done. 
\end{proof}

Due to Lemma ~\ref{lemma: x_1^1,x_2^2,x_3^1}, it is sufficient to consider any arbitrary sequences $x_1$, $x_2$ and $x_3$ in 
$\mathbb{F}_2^{n}$ such that $x_1^1=1$, $x_2^1=0$ and $x_3^1=0$ or $x_1^1=0$, $x_2^1=1$ and $x_3^1=1$.
Due to the symmetry it suffices to deal with the case where $x_1^1=1$, $x_2^1=0$ and $x_3^1=0$. We need a few lemmas.

\begin{lemma} \label{lemma: x_1^2=x_2^2=x_3^2}
Assume that $t \geq 1$ and $n \geq t+2$. Let $x_1$,$x_2$ and $x_3$ be three arbitrary sequences in $\mathbb{F}_2^n$ such that $x_1^1=1$, $x_2^1=0$ and $x_3^1=0$. If $x_1^2=x_2^2=x_3^2$, then 
$$|\mathcal{D}_t(x_1) \cap \mathcal{D}_t(x_2) \cap \mathcal{D}_t(x_3)|\leq N_3(n,t).$$
\end{lemma}
\begin{proof}
Case 1: Assume that $x_1^2=x_2^2=x_3^2=0$.The we have 
$x_1=10x_1^{\prime}$, $x_2=00x_2^{\prime}$ and $x_3=00x_3^{\prime}$.
Note that $x_2^{\prime} \neq x_3^{\prime}$. Suppose $x_2^k$ is the first occurrence of the symbol 1 in $x_2$ and the symbol 1 appears in $x_2$ not after it appears in $x_3$. If $x_2^k=x_3^k=1$ we have 
\begin{align*}
|\chi^{1}|& \leq |1 \circ \mathcal{D}_{t-(k-1)}(x_2^{k+1},...,x_2^n) \cap \mathcal{D}_{t-(k-1)}(x_3^{k+1},...,x_3^n)|\\
&\leq \max_{x \neq y, x,y \in \mathbb{F}_2^{n-k}}|\mathcal{D}_{t-(k-1)}(x)\cap \mathcal{D}_{t-(k-1)}(y)|\\
& \leq 2D(n-2-k,t-k)\\
& \leq 2D(n-5,t-3).
\end{align*}

If $x_3^k=0$ we have 
\begin{align*}
|\chi^{1}|& \leq |1 \circ \mathcal{D}_{t-k}(x_3^{k+2},...,x_3^n)|\\
&\leq D(n-k-1,t-k)\\
& \leq D(n-4,t-3)\\
& \leq 2D(n-5,t-3).
\end{align*}
We also have that 
\begin{align*}
|\chi^{0}|& \leq |0 \circ \mathcal{D}_{t-1}(x_1^{\prime})|\\
& \leq D(n-2,t-1).
\end{align*}
Since $D(n-2,t-1)=D(n-3,t-1)+D(n-4,t-2)$ and $D(n-5,t-3) \leq D(n-4,t-2)$, we have 
$$|\chi^{1}|+|\chi^{0}| \leq N_3(n,t).$$

Case 2: Assume that $x_1^2=x_2^2=x_3^2=1$. Suppose $x_1^k$  is the first occurrence of the symbol 0 in $x_1$. We have that 
\begin{align*}
|\chi^{0}|& \leq |0 \circ \mathcal{D}_{t-(k-1)}({x_1^{k+1},...,x_1^n})|\\
& \leq D(n-k,t-k+1)\\
& \leq D(n-3,t-2).
\end{align*}
We also have 
\begin{align*}
|\chi^{1}|& \leq |1 \circ \mathcal{D}_{t-1}(\mathbf{x_2}^{\prime}) \cap \mathcal{D}_{t-1}(\mathbf{x_3}^{\prime})|\\
& \leq 2D(n-4,t-2).
\end{align*} 
Since 
\begin{align*}
D(n-3,t-2) &=D(n-4,t-2)+D(n-5,t-3)\\
& \leq D(n-3,t-1)+D(n-5,t-3),
\end{align*} 
we have 
$$|\chi^{1}|+|\chi^{0}| \leq N_3(n,t). \qedhere$$
\end{proof}

Now we will show that it is sufficient to consider $x_1$,$x_2$ and $x_3$ three arbitrary sequences in $\mathbb{F}_2^n$ such that $x_1^1=1$, $x_2^1=0$, $x_3^1=0$,$x_1^2=0$, $x_2^2=1$ and $x_3^2=1$.

\begin{lemma} \label{x_1^1=1,x_2^1=0,x_3^1=0}
Assume $t \geq 1$ and $n \geq t+2$. Let $x_1$,$x_2$ and $x_3$ be three arbitrary sequences in $\mathbb{F}_2^n$ such that $x_1^1=1$, $x_2^1=0$, $x_3^1=0$.

(a)$x_1^2=1$, $x_2^2=0$ and $x_3^2=0$.

(b)$x_1^2=1$, $x_2^2=1$ and $x_3^2=0$. 

(c)$x_1^2=0$, $x_2^2=1$ and $x_3^2=0$. 

Then under (a),or (b), or (c)
$$|\mathcal{D}_t(x_1) \cap \mathcal{D}_t(x_2) \cap \mathcal{D}_t(x_3)|\leq N_3(n,t).$$
\end{lemma}
\begin{proof}
Let $\chi=\mathcal{D}_t(x_1) \cap \mathcal{D}_t(x_2) \cap \mathcal{D}_t(x_3).$

Under (a):

Let $x_1^k$ be the first occurrence of symbol 0 in $x_1$, we have
\begin{align*}
|\chi^{0}|& \leq |0 \circ \mathcal{D}_{t-(k-1)}(x_1^{k+1},...,x_1^n)|\\
&\leq D(n-k,t-k+1)\\
& \leq D(n-3,t-2).
\end{align*}
Similarly, let $x_2^k$ be the first occurrence of symbol 1 in $x_2$, we have
\begin{align*}
|\chi^{1}|& \leq |0 \circ \mathcal{D}_{t-(k-1)}(x_2^{k+1},...,x_2^n)|\\
&\leq D(n-k,t-k+1)\\
& \leq D(n-3,t-2).
\end{align*}
Therefore, we have
$$|\chi^{1}|+|\chi^{0}| \leq N_3(n,t).$$

Under (b):

Let $x_1^k$ be the first occurrence of symbol 0 in $x_1$ we have
\begin{align*}
|\chi^{0}|& \leq |0 \circ \mathcal{D}_{t-(k-1)}(x_1^{k+1},...,x_1^n)|\\
&\leq D(n-k,t-k+1)\\
& \leq D(n-3,t-2).
\end{align*}
Similarly, let $x_3^k$ be the first occurrence of symbol 1 in $x_3$, we have
\begin{align*}
|\chi^{1}|& \leq |0 \circ \mathcal{D}_{t-(k-1)}(x_3^{k+1},...,x_3^n)|\\
&\leq D(n-k,t-k+1)\\
& \leq D(n-3,t-2).
\end{align*}
Therefore, we have
$$|\chi^{1}|+|\chi^{0}| \leq N_3(n,t).$$

Under(c):

If $x_3^k$ be the first occurrence of symbol 1 in $x_3$, we have 
\begin{align*}
|\chi^{1}|& \leq |0 \circ \mathcal{D}_{t-(k-1)}(x_3^{k+1},...,x_3^n)|\\
&\leq D(n-k,t-k+1)\\
& \leq D(n-3,t-2).
\end{align*}

\begin{align*}
|\chi^{0}|& \leq |0 \circ \mathcal{D}_{(t-1)}(x_1^{3},...,x_1^n)|\\
& \leq D(n-2,t-1).
\end{align*}
Therefore, we have
$$|\chi^{1}|+|\chi^{0}| \leq N_3(n,t). \qedhere$$
\end{proof}

\begin{lemma} \label{x_1^{12}}
Assume that $t \geq 1$ and $n \geq t+2$. Let $x_1$,$x_2$ and $x_3$ be three arbitrary sequences in $\mathbb{F}_2^n$ such that $x_1^1=1$, $x_2^1=0$ and $x_3^1=0$ and $x_1^2=0$, $x_2^2=1$ and $x_3^1=1$  Then we have 
$$|\mathcal{D}_t(x_1) \cap \mathcal{D}_t(x_2) \cap \mathcal{D}_t(x_3)|\leq N_3(n,t).$$
\end{lemma}
\begin{proof}
Let $\chi=\mathcal{D}_t(x_1) \cap \mathcal{D}_t(x_2) \cap \mathcal{D}_t(x_3).$ We have 
\begin{align*}
|\chi^{1}|& \leq |1 \circ \mathcal{D}_{t-1}(x_2^{3},...,x_2^n) \cap \mathcal{D}_{t-1}(x_3^{3},...,x_3^n)|\\
&\leq \max_{x \neq y,x,y \in \mathbb{F}_2^{n-2}}|\mathcal{D}_{t-1}(x)\cap \mathcal{D}_{t-1}(y)|\\
& \leq 2D(n-4,t-2).
\end{align*}

Also 
\begin{align*}
|\chi^{0}|& \leq |0 \circ \mathcal{D}_{(t-1)}(x_1^{3},...,x_1^n)|\\
& \leq D(n-2,t-1).
\end{align*}
Since $2D(n-4,t-2)+D(n-2,t-1)=N_3(n,t)$, the proof is done. 
\end{proof}
By Lemmas \ref{lemma: x_1^1,x_2^2,x_3^1}, \ref{lemma: x_1^2=x_2^2=x_3^2}, \ref{x_1^1=1,x_2^1=0,x_3^1=0}, and \ref{x_1^{12}}, the induction step is finished and we have the following theorem.
\begin{theorem}
For $t \geq 1$ and $n \geq t+2$, we have
$$N(n,3,t) \leq N_3(n,t)$$
\end{theorem}

\section{The Intersection of $\ell \geq 4$ Deletion Balls}
In this section we will prove that 
$$N(n,\ell,t)=N_{\ell}(n,t):=\sum_{i=1}^{\ell-2}D(n-2i,t-i)+2D(n-2(\ell-1),t-(\ell-1)).$$

\subsection{The Lower Bound}
We will show that $N_{\ell}(n,t)$ is a lower bound for $N(n,\ell,t)$ by showing that $x_1=10a_{n-2}$, $x_2=01a_{n-2}$, $x_j=\underbrace{0101\cdots01}_{2(j-1)}a_{n-2(j-1)}$  for $j \in \{2,\cdots,\ell\}$ satisfy 
$$|\cap_{i=1}^{\ell}\mathcal{D}_t(x_i)|=N_{\ell}(n,t).$$

\begin{theorem}
For $ \ell \geq 4$, $t \geq 1$, and $n \geq t+\ell-1$,
$$N(n,\ell,t) \geq \sum_{i=1}^{\ell-2}D(n-2i,t-i)+2D(n-2(\ell-1),t-(\ell-1))$$
\end{theorem}

\begin{proof}
Let $x_1=10a_{n-2}$, $x_2=01a_{n-2}$, $x_j=\underbrace{0101\cdots01}_{2(j-1)}a_{n-2(j-1)}$  for $j \in \{2,\cdots,\ell\}$ and let $\chi=\cap_{i=1}^{\ell}\mathcal{D}_t(x_i)$. Observe that 
$$(0 \circ \mathcal{D}_{(t-1)}(x_1^3,...,x_1^n)) \cap (0 \circ \mathcal{D}_{t}(x_2^2,...,x_2^n)) \cap \cdots \cap (0 \circ \mathcal{D}_{t}(x_{\ell}^2,...,x_{\ell}^n))=0 \circ \mathcal{D}_{(t-1)}(x_1^3,...,x_1^n)$$
and 
\begin{align*}
 (1 \circ \mathcal{D}_{t}(x_1^2,...,x_1^n)) \cap (1 \circ \mathcal{D}_{t-1}(x_2^3,...,x_2^n)) \cap \cdots \cap (1 \circ \mathcal{D}_{t-1}(x_{\ell}^3,...,x_{\ell}^n))\\
=(1 \circ \mathcal{D}_{t-1}(x_2^3,...,x_2^n)) \cap (1 \circ \mathcal{D}_{t-1}(x_3^3,...,x_3^n)) \cdots \cap (1 \circ \mathcal{D}_{t-1}(x_{\ell}^3,...,x_{\ell}^n))
\end{align*}

By Lemma 1 and Lemma 2, we have 
\begin{align*}
|\chi^{0}|& = |(0 \circ \mathcal{D}_{(t-1)}(x_1^3,...,x_1^n)) \cap (0 \circ \mathcal{D}_{t}(x_2^2,...,x_2^n)) \cap \cdots \cap (0 \circ \mathcal{D}_{t}(x_{\ell}^2,...,x_{\ell}^n))|\\
&=|(0 \circ \mathcal{D}_{t-1}(x_1^3,...,x_1^n))|\\
& = D(n-2,t-1)
\end{align*}
\begin{align*}
|\chi^{1}|& = |(1 \circ \mathcal{D}_{t}(x_1^2,...,x_1^n)) \cap (1 \circ \mathcal{D}_{t-1}(x_2^3,...,x_2^n)) \cap (1 \circ \mathcal{D}_{t-1}(x_3^3,...,x_3^n)) \cap \cdots \cap (1 \circ \mathcal{D}_{t-1}(x_{\ell}^3,...,x_{\ell}^n))|\\
&=|(1 \circ \mathcal{D}_{t-1}(x_2^3,...,x_2^n)) \cap (1 \circ \mathcal{D}_{t-1}(x_3^3,...,x_3^n)) \cdots \cap (1 \circ \mathcal{D}_{t-1}(x_{\ell}^3,...,x_{\ell}^n))|\\
& = N_{\ell-1}(n-2,t-1)
\end{align*}
Therefore,
\begin{align*}
|\chi|& =|\chi^{0}|+|\chi^{1}|\\
& = D(n-2,t-1)+N_{\ell-1}(n-2,t-1). \qedhere
\end{align*}
\end{proof}

Now we have the following recursive relations. 
$$N_2(n,t)=2D(n-2,t-1)$$
$$N_3(n,t)=D(n-2,t-1)+2D(n-4,t-2)$$
$$N_4(n,t)=D(n-2,t-1)+N_3(n-2,t-1)$$
...
$$N_{\ell}(n,t)=D(n-2,t-1)+N_{\ell-1}(n-2,t-1).$$

Therefore, 
$$N_{\ell}(n,t)=\sum_{i=1}^{\ell-2}D(n-2i,t-i)+2D(n-2(\ell-1),t-(\ell-1)).$$

\subsection{The Upper Bound}

 We will show that $N_{\ell}(n,t)$ is an upper bound for $N(n,\ell,t)$ by induction on $\ell$, $n$, $t$. Let us first address the base case. 

The base case is: 
$n=t+\ell-1$, $t \geq 1$, and $\ell \geq 3$ or

$t=1$, $\ell \geq 3$, and $n \geq t+\ell-1$ or 

$\ell=3$ , $t \geq 1$,  and $n \geq t+\ell-1$.

$n=t+\ell-1$, $t \geq 1$, and $\ell \geq 3$, it is easy to calculate that $N_{\ell}(t+\ell-1,t)=2^{\ell-1}$. After deleting $t$ symbols from each sequence of length $t+\ell-1$ there are only $\ell-1$ symbols left for each sequence. Since $N_{\ell}(n,t)=2^{\ell-1}$, $N(n,\ell,t) \leq N_{\ell}(n,t)$. 

For $t=1$, $\ell \geq 3$, and $n \geq t+\ell-1$, we have $N_\ell(n,1)=1$. Since $N(n,3,1)=1$, we have $N(n,\ell,1) \leq 1=N_{\ell}(n,1)$. 

For $\ell=3$ , $t \geq 1$,  and $n \geq t+\ell-1$, in the previous section it has been proven. 

Now we will move onto the induction step. Assume that $N(n_0,\ell_0,t_0) \leq N_{\ell_0}(n_0,t_0)$ is true for all $n_0 \geq t_0+2$, $\ell_0 \geq 3$, and $t_0\geq 1$ such that $n_0+t_0+\ell_0 < n+t+\ell$. We will need a few lemmas to complete this step.

Let us first show the following two lemmas which will be used in our analysis several times.
\begin{lemma} \label{lemma: D(x-1,y-1)}
For any integers $x,y$ and any $m\ge 2$,
$$ D(x-1,y-1)\ \le\ N_{m}(x,y).$$

\end{lemma}

\begin{proof}
We argue by induction on $m\ge 2$.

\medskip
For the base case let $m=2$. 
Here $N_2(x,y)=2D(x-2,y-1)$. 

We have 
$$
D(x-1,y-1)=D(x-2,y-1)+D(x-3,y-2).
$$
Moreover,
$$
D(x-3,y-2)\ \le\ D(x-2,y-1),
$$
Thus
$$
D(x-1,y-1)\ \le\ 2D(x-2,y-1)\ =\ N_2(x,y).
$$

For the induction step assume the claim holds for some $m-1\ge 2$, i.e.,
$$
D(u-1,v-1)\ \le\ N_{m-1}(u,v)\qquad\text{for all }u,v\in\mathbb{Z}.
$$
We show it for $m$. Using the same recursion,
$$
D(x-1,y-1)=D(x-2,y-1)+D(x-3,y-2).
$$
On the other hand, by the definition of $N_m$,
$$
N_m(x,y)=D(x-2,y-1)+N_{m-1}(x-2,y-1).
$$

Applying the induction hypothesis gives
$$
D(x-3,y-2)\ \le\ N_{m-1}(x-2,y-1).
$$
Therefore
$$
D(x-1,y-1)
\ \le\ D(x-2,y-1)+N_{m-1}(x-2,y-1)
\ =\ N_m(x,y),
$$
completing the induction.
\end{proof}

\begin{lemma} \label{lemma: N_m,N_L}
For any $m$ and any $L\ge m$,
$$
N_m(u-1,v-1)\ \le\ N_L(u,v).
$$
\end{lemma}
\begin{proof}
Let $L=m$. Since $D(n,t) \geq D(n-1,t-1)$, we have 
$$
N_m(u-1,v-1) \leq N_L(u,v). 
$$
Now assume that $L>m$.  We will bound $N_m(u-1,v-1)$ from above termwise.

For $1\le i\le m-2$, we have:
$$
D(u-1-2i,\,v-1-i)\ \le\ D(u-2i,\,v-i).
$$
Since $L> m$, the indices $i=1,\dots,m-2$ all lie in the index
range $1\le i\le L-2$ of $N_L(u,v)$. Thus each of these terms is dominated by
a corresponding term of $N_L(u,v)$.

Let $a=u-1-2(m-1)$ and $b=v-1-(m-1)$. Then
$$
2D(a,b)=2D\bigl(u-1-2(m-1),\,v-1-(m-1)\bigr).
$$

If $L>m$, then $m-1\le L-2$. We have 
$$
2D(a,b)
\le D\bigl(u-2(m-1)+1,\,v-(m-1)\bigr)
\le D\bigl(u-2(m-1),\,v-(m-1)\bigr).
$$

\medskip
\noindent
Therefore, we have 
$$
N_m(u-1,v-1)\ \le\ N_L(u,v). \qedhere
$$
\end{proof}

Now let us assume that $n \geq t+\ell-1$, $t \geq 1$ and $\ell \geq 4$. 

\begin{lemma} \label{lemma: l_1 10,l_2 01}
Suppose that among the vectors $x_1, x_2, \dots, x_{\ell}$, 
$\ell_1$ of them begin with $10$ and $\ell_2$ of them begin with $01$, 
where $\ell_1 \geq 2$ and $\ell_2 \geq 2$. Then
$$|\mathcal{D}_t(\mathbf{x}_1) \cap \mathcal{D}_t(\mathbf{x}_2) \cap ... \cap \mathcal{D}_t(\mathbf{x}_{\ell}) |\leq N_{\ell}(n,t).$$
\end{lemma}
\begin{proof}
Let $\chi= \cap_{i=1}^{\ell} \mathcal{D}_{t}(x_i)$ and then 
$\chi^0=0 \circ \cap_{i=1}^{\ell_1}\mathcal{D}_{t-1}(x_i^3x_i^4\cdots x_i^n)$.
$$|\chi^0| \leq \max_{y_i \in \{0,1\}^{n-2}} \cap_{i=1}^{\ell_1} \mathcal{D}_{t-1}(y_i) \leq N_{\ell_1}(n-2,t-1).$$

Similarly, we have 
$$|\chi^1| \leq \max_{y_i \in \{0,1\}^{n-2}} \cap_{i=1}^{\ell_2} \mathcal{D}_{t-1}(y_i) \leq N_{\ell_2}(n-2,t-1).$$

Let $\ell,\ell_1,\ell_2 \ge 2$ with $\ell_1+\ell_2=\ell$ and $\ell\ge 4$. We will show that 
$$
N_{\ell}(n,t) \ge N_{\ell_1}(n-2,t-1)+N_{\ell_2}(n-2,t-1).
$$

We first note the identity
\begin{equation}\label{eq:D-recursion}
D(n,t)=D(n-1,t) + D(n-2,t-1).
\end{equation}
Since
$$
D(n-1,t)-D(n-2,t-1)
=\binom{n-t-1}{t}\ge 0,
$$
we obtain from \eqref{eq:D-recursion} that
\begin{equation}\label{eq:D-ineq}
D(n,t)\ge 2D(n-2,t-1).
\end{equation}

Applying \eqref{eq:D-ineq} to each term in the definition of $N_\ell(n,t)$ gives
\begin{align}
N_{\ell}(n,t)
&=\sum_{i=1}^{\ell-2} D(n-2i,t-i)
\;+\;2D(n-2(\ell-1),t-(\ell-1))\\
&\ge 2\sum_{i=1}^{\ell-2} D(n-2-2i,t-1-i)
\;+\;2D(n-2-2(\ell-1),t-\ell).\label{eq:N-ineq}
\end{align}

Now we expand $N_{\ell_1}(n-2,t-1)$ and $N_{\ell_2}(n-2,t-1)$:
\begin{align*}
N_{\ell_1}(n-2,t-1)
&=\sum_{i=1}^{\ell_1-2} D(n-2-2i,t-1-i)
\;+\;2D(n-2-2(\ell_1-1),t-\ell_1),\\
N_{\ell_2}(n-2,t-1)
&=\sum_{j=1}^{\ell_2-2} D(n-2-2j,t-1-j)
\;+\;2D(n-2-2(\ell_2-1),t-\ell_2).
\end{align*}

Summing, we obtain
\begin{align}
&N_{\ell_1}(n-2,t-1)+N_{\ell_2}(n-2,t-1)\\
&=\sum_{i=1}^{\ell_1-2} D(n-2-2i,t-1-i)
+\sum_{j=1}^{\ell_2-2} D(n-2-2j,t-1-j)\\
&+2D(n-2-2(\ell_1-1),t-\ell_1)
+2D(n-2-2(\ell_2-1),t-\ell_2).\label{eq:sum-l1l2}
\end{align}

 Also we have 
$$D(n-2(\ell_1-1),\,t-(\ell_1-1)) \geq 2D(n-2-2(\ell_1-1),\,t-\ell_1), $$

and 
$$D(n-2(\ell_2-1),\,t-(\ell_2-1)) \geq  2D(n-2-2(\ell_2-1),\,t-\ell_2).$$

Since $\ell_1-1,\ell_2-1\le \ell-2$ when $\ell_1,\ell_2\ge 2$, every term in \eqref{eq:sum-l1l2} appears among the summands in \eqref{eq:N-ineq} with coefficient at most $2$. The additional term $2D(n-2-2(\ell-1),t-\ell)\ge 0 $ in \eqref{eq:N-ineq} only increases the right-hand side of that inequality.

Therefore, comparing \eqref{eq:N-ineq} and \eqref{eq:sum-l1l2} yields
$$
N_{\ell}(n,t)\ge N_{\ell_1}(n-2,t-1)+N_{\ell_2}(n-2,t-1),$$
which completes the proof.
\end{proof}

\begin{lemma} \label{lemma:x_1=x_2=a}
 Let $x_1$,$x_2$,$\cdots$ and $x_{\ell}$ be $\ell$ arbitrary sequences in $\mathbb{F}_2^n$ such that $x_i \neq x_j$ for $i \neq j$ and $a=x_1^1=x_2^1=\cdots =x_{\ell}^1$. Then 
$$|\cap_{i}^{\ell}\mathcal{D}_t(x_i) |\leq N_{\ell}(n,t).$$
\end{lemma}
\begin{proof}
    Our proof follows the proof of Theorem 8 in \cite{Gabrys}. Let $\chi=\mathcal{D}_t(x_1) \cap \mathcal{D}_t(x_2) \cap \cdots \cap \mathcal{D}_t(x_{\ell})$. We have 
\begin{align*}
|\chi^{a}|& =|a \circ \mathcal{D}_{t}(x_1^2,...,x_1^n)\cap \mathcal{D}_{t}(x_2^2,...,x_2^n) \cap \cdots \cap \mathcal{D}_{t}(x_{\ell}^2,...,x_{\ell}^n)|\\
&\leq \max_{x_1 \neq x_2 \neq \cdots x_{\ell}; x_1,x_2,\cdots, x_{\ell} \in \mathbb{F}_2^{n-1}}|\mathcal{D}_t(x_1) \cap \mathcal{D}_{t}(x_2)\cap \cdots \cap \mathcal{D}_{t}(x_{\ell})|\\
& \leq N_{\ell}(n-1,t)
\end{align*}

Suppose $x_1^k$ is the first occurrence of the symbol $\bar{a}$ in $x_1$ and the symbol $\bar{a}$ appears in $x_1$ not after it appears in $x_2$, $x_3$, $\cdots$ and $x_\ell$.

If $x_1^k=x_2^k=\cdots =x_{\ell}^k=\bar{a}$ we have 
\begin{align*}
|\chi^{\bar{a}}|& =|\bar{a} \circ \mathcal{D}_{t-(k-1)}(x_1^{k+1},...x_1^n)\cap \mathcal{D}_{t-(k-1)}(x_2^{k+1},...,x_2^n) \cap \cdots \cap \mathcal{D}_{t-(k-1)}(x_{\ell}^{k+1},...,x_{\ell}^n)|\\
&\leq \max_{\mathbf{x_1} \neq \mathbf{x_2} \cdots \neq \mathbf{x_{\ell}};x_1, x_2,\cdots,x_{\ell}\in \mathbb{F}_2^{n-k}}|\mathcal{D}_{t-(k-1)}(x_1) \cap \mathcal{D}_{t-(k-1)}(x_2)\cap \cdots \cap \mathcal{D}_{t-(k-1)}(x_{\ell})|\\
& \leq N_{\ell}(n-k,t-k+1)\\
& \leq N_{\ell}(n-2,t-1)
\end{align*}
If one of $x_2^k$, $x_3^k$ $\cdots$ or $x_{\ell}^k$ is equal to $a$, say $x_2^k=a$ then we have 
\begin{align}
|\chi^{\bar{a}}|& \leq |\bar{a} \circ  \mathcal{D}_{t-k}(x_2^{k+2},...x_2^n)|\\
&\leq D(n-k-1,t-k)\\
& \leq D(n-3,t-2)\\
& \leq N_{\ell}(n-2,t-1) \label{eq: D(n-3,t-2) < N}
\end{align}
Let us show \eqref{eq: D(n-3,t-2) < N}.

By using \eqref{eq:D-recursion}, we prove by induction that for any integer $k\ge1$,
\begin{equation}\label{eq:telescope}
D(n-3,t-2)
=
\sum_{i=1}^k D(n-2-2i,\ t-1-i)
+
D(n-(2k+3),\ t-(k+2)).
\end{equation}

Let $k=\ell-2$. Then \eqref{eq:telescope} becomes
\begin{equation}\label{eq:k=l-2}
D(n-3,t-2)
=
\sum_{i=1}^{\ell-2} D(n-2-2i,\ t-1-i)
+
D\bigl(n-(2\ell-1),\ t-\ell).
\end{equation}

Note that from \begin{equation}
D(n,t)=D(n-1,t) + D(n-2,t-1),
\end{equation} 
and 
$$
D(n-1,t)-D(n-2,t-1)
=\binom{n-t-1}{t}\ge 0,$$

we have $D(n,t) \leq 2D(n-1,t)$. Thus  
$$
D(n-(2\ell-1),\,t-\ell)
\le
2\,D(n-2\ell,\,t-\ell)
=
2\,D(n-2-2(\ell-1),\,t-\ell).
$$

Substituting this estimate into \eqref{eq:k=l-2} yields
$$
D(n-3,t-2)
\le
\sum_{i=1}^{\ell-2} D(n-2-2i,\ t-1-i)
+
2\,D(n-2-2(\ell-1),\,t-\ell)
=
N_{\ell}(n-2,t-1).$$

Since $|\chi^a|+|\chi^{\bar{a}}|\leq N_{\ell}(n-1,t)+N_{\ell}(n-2,t-1)=N_{\ell}(n,t)$, the proof is completed.
\end{proof}

\begin{lemma}
Suppose that among the vectors $x_1, x_2, \dots, x_{\ell}$, 
$\ell_1 \geq 1$ of them begin with $11$ and $\ell_3 \geq 1$ of them begin with $00$. Then
$$|\mathcal{D}_t(x_1) \cap \mathcal{D}_t(x_2) \cap ... \cap \mathcal{D}_t(x_{\ell}) |\leq N_{\ell}(n,t).$$
\end{lemma}
\begin{proof}
Assume that the first two coordinates of $x_1$, $x_2$,$\cdots$, $x_{\ell_1}$ are $11$ and the first two coordinates of $x_{\ell_1+1}$, $x_{\ell_1+2}$, $\cdots$, $x_{\ell_1+\ell_3}$ are $00$. Let $\chi= \cap_{i=1}^{\ell} \mathcal{D}_{t}(x_i)$

Case 1: $\ell_1 \geq 2$ and $\ell_3 \geq 2$.

Then we have 
$\chi^0 \subset 0 \circ \max_{y_i: y_i \in \{0,1\}^{n-3}}\cap_{i=1}^{\ell_1}\mathcal{D}_{t-2}(y_i)$.
$$|\chi^0| \leq \max_{y_i \in \{0,1\}^{n-3}, y_i \neq y_j} \cap_{i=1}^{\ell_1} \mathcal{D}_{t-2}(y_i) \leq N_{\ell_1}(n-3,t-2).$$
Similarly, we have 
$\chi^1 \subset 1 \circ \max_{y_i: y_i \in \{0,1\}^{n-3}}\cap_{i=1}^{\ell_3}\mathcal{D}_{t-2}(y_i)$.
$$|\chi^0| \leq \max_{y_i \in \{0,1\}^{n-3}, y_i \neq y_j} \cap_{i=1}^{\ell_3} \mathcal{D}_{t-2}(y_i) \leq N_{\ell_3}(n-3,t-2).$$

By the proof of Lemma ~\ref{lemma: l_1 10,l_2 01}, we have 
 $$N_{\ell_1}(n-2,t-1)+N_{\ell_3}(n-2,t-1)
 \le\ N_{\ell}(n,t). $$
Therefore, 
\[
N_{\ell_1}(n-3,t-2)+N_{\ell_3}(n-3,t-2)
\ \le\ N_{\ell_1}(n-2,t-1)+N_{\ell_3}(n-2,t-1)
\ \le\ N_{\ell}(n,t).
\]

Case 2: $\ell_1=1$ and $\ell_3 \geq 2$.
$\chi^0 \subset 0 \circ \max_{y: y \in \{0,1\}^{n-3}}\mathcal{D}_{t-2}(y)$.
$$|\chi^0| \leq \max_{y_i \in \{0,1\}^{n-3}}  \mathcal{D}_{t-2}(y_i) \leq D(n-3,t-2).$$

Similarly, we have 
$\chi^1 \subset 1 \circ \max_{y_i: y_i \in \{0,1\}^{n-3}}\cap_{i=1}^{\ell_3}\mathcal{D}_{t-2}(y_i)$.
$$|\chi^0| \leq \max_{y_i \in \{0,1\}^{n-3}, y_i \neq y_j} \cap_{i=1}^{\ell_3} \mathcal{D}_{t-2}(y_i) \leq N_{\ell_3}(n-3,t-2).$$

We claim that  $D(n-3,t-2)+ N_{\ell_3}(n-3,t-2) \leq N_{\ell}(n,t) $. Let us prove this claim. 

We have 
\begin{equation*}
N_{\ell}(n,t)
= D(n-2,t-1)\;+\;N_{\ell-1}(n-2,t-1).
\end{equation*}
and 
\[
D(n-3,t-2)\ \le\ D(n-2,t-1).
\]

Apply Lemma ~\ref{lemma: N_m,N_L} with $m=\ell_3$, $L=\ell-1$, and $(u,v)=(n-2,t-1)$ to get
\[
N_{\ell_3}(n-3,t-2)\ \le\ N_{\ell-1}(n-2,t-1),
\]
since $\ell_3\le \ell-1$ by hypothesis.

Therefore,
\[
D(n-3,t-2)+N_{\ell_3}(n-3,t-2)
\ \le\ D(n-2,t-1)+N_{\ell-1}(n-2,t-1)
\ =\ N_{\ell}(n,t).
\]

Case 3: $\ell_1=1$ and $\ell_3=1$.
We have $\chi^0 \subset 0 \circ \max_{y: y \in \{0,1\}^{n-3}}\mathcal{D}_{t-2}(y)$.
$$|\chi^0| \leq \max_{y_i \in \{0,1\}^{n-3}}  \mathcal{D}_{t-2}(y_i) \leq D(n-3,t-2),$$
and 
$\chi^1 \subset 1 \circ \max_{y: y \in \{0,1\}^{n-3}}\mathcal{D}_{t-2}(y)$.
$$|\chi^0| \leq \max_{y_i \in \{0,1\}^{n-3}}  \mathcal{D}_{t-2}(y_i) \leq D(n-3,t-2).$$

By Lemma ~\ref{lemma: D(x-1,y-1)}, we have 
$$2D(n-3,t-2) \leq 2N_2(n-2,t-1) \leq N_{\ell}(n,t),$$
since $\ell \geq 4$.
\end{proof}

Now we can assume none of the $\ell$ vectors start with $00$.

\begin{lemma}
   Suppose that among the vectors $x_1, x_2, \dots, x_{\ell}$, 
$\ell_1 \geq 1$ of them begin with $11$, $\ell_2 \geq 1$ of them begin with $01$, 
and $\ell_4$ of them begin with $10$. Then
$$|\mathcal{D}_t(x_1) \cap \mathcal{D}_t(x_2) \cap ... \cap \mathcal{D}_t(x_{\ell}) |\leq N_{\ell}(n,t).$$
\end{lemma}
\begin{proof}
Let $\chi= \cap_{i=1}^{\ell} \mathcal{D}_{t}(x_i)$

Case 1: $\ell_1 \geq 1$ and $\ell_2 \geq 2$. 

Assume that the first two coordinates of $x_1$,$x_2$,$\cdots$, $x_{\ell_1}$ are $11$ and the first two coordinates of $x_{\ell_1+1}$,$x_{\ell_1+2}$,$\cdots$, $x_{\ell_1+\ell_2}$ are $01$. Then we have $\chi^0 \subset 0 \circ \max_{y: y \in \{0,1\}^{n-3}}\mathcal{D}_{t-2}(y)$.
$$|\chi^0| \leq \max_{y_i \in \{0,1\}^{n-3}}  \mathcal{D}_{t-2}(y_i) \leq D(n-3,t-2).$$

Similarly, $\chi^1 \subset 1 \circ \max_{y_1,y_2 \in \{0,1\}^{n-2}}\mathcal{D}_{t-1}(y_1) \cap \mathcal{D}_{t-1}(y_2)$.
$$|\chi^1| \leq \max_{y_1,y_2 \in \{0,1\}^{n-2}}  \mathcal{D}_{t-1}(y_1) \cap \mathcal{D}_{t-1}(y_2) \leq 2D(n-4,t-2).$$

We have 
\begin{align}
N_{\ell}(n,t)
&= D(n-2,t-1) + N_{\ell-1}(n-2,t-1). \label{eq:decomp}
\end{align}

Note that we have 
\[
D(n-2,t-1)\ \ge\ 2\,D(n-4,t-2).
\]
By Lemma ~\ref{lemma: D(x-1,y-1)}:
\[
N_{\ell-1}(n-2,t-1)\ \ge\ D\bigl((n-2)-1,(t-1)-1\bigr)\ =\ D(n-3,t-2).
\]
Combining these two bounds in \eqref{eq:decomp} yields
\[
N_{\ell}(n,t)\ \ge\ 2\,D(n-4,t-2)\;+\;D(n-3,t-2),
\]
as claimed.

Case 2: $\ell_1 \geq 2$ and $\ell_2=1$.

We have $\chi^1 \subset 1 \circ \max_{y \in \{0,1\}^{n-2}}\mathcal{D}_{t-1}(y)$.
$$|\chi^1| \leq \max_{y \in \{0,1\}^{n-2}}  \mathcal{D}_{t-1}(y) \leq D(n-2,t-1).$$

$\chi^0 \subset 0 \circ \max_{y_i: y_i \in \{0,1\}^{n-3}}\cap_{i=1}^{\ell_1}\mathcal{D}_{t-2}(y_i)$.
$$|\chi^0| \leq \max_{y_i \in \{0,1\}^{n-3}, y_i \neq y_j} \cap_{i=1}^{\ell_1} \mathcal{D}_{t-2}(y_i) \leq N_{\ell_1}(n-3,t-2).$$

By lemma ~\ref{lemma: D(x-1,y-1)}, we have 
$$D(n-2,t-1) \leq N_\ell(n-1,t).$$
Additionally we have $N_{\ell_1}(n-3,t-2) \leq N_\ell(n-2,t-1)$ by Lemma ~\ref{lemma: N_m,N_L}. 

Finally, $N_\ell(n-2,t-1)+N_\ell(n-1,t)=N_{\ell}(n,t)$

Case 3: $\ell_1= 1$, $\ell_2=1$ and $\ell_4=\ell-2$.

We have $\chi^1 \subset 1 \circ \max_{y \in \{0,1\}^{n-2}}\mathcal{D}_{t-1}(y)$.
$$|\chi^1| \leq \max_{y \in \{0,1\}^{n-2}}  \mathcal{D}_{t-1}(y) \leq D(n-2,t-1).$$

Also we have since $k \geq 3$
\begin{align*}
\chi^0 &  \subset 0 \circ \max_{x_i \in \{0,1\}^{n-2}, y \in \{0,1\}^{n-k}} \cap_{i=1}^{\ell_4} \mathcal{D}_{t-1}(x_i) \cap \mathcal{D}_{t-k+1}(y)\\
& \subset 0 \circ \max_{y \in \{0,1\}^{n-3}} \mathcal{D}_{t-2}(y).
\end{align*}

we have $|\chi^{0}| \leq D(n-3,t-2) \leq  N_{\ell}(n-2,t-1).$
By Lemma ~\ref{lemma: D(x-1,y-1)}, we have $D(n-2,t-1) \leq N_{\ell}(n-1,t)$. 
\end{proof}
Suppose that among the vectors $x_1, x_2, \dots, x_{\ell}$, one begins with $10$ 
and another begins with $11$. In this case, by symmetry, the situation is equivalent 
to that considered in the preceding lemma.

Now the induction step is finished. we have the following theorem. 

\begin{theorem}
For $t \geq 1$, $\ell \geq 3$, and $n \geq t+\ell-1$, we have
$$N(n,\ell,t) \leq N_{\ell}(n,t).$$
\end{theorem}

\section{Conclusion and Open Problems}
In this paper we derive the maximum size of the intersection of the $t$-deletion balls centered at $x_1,x_2,\cdots,x_{\ell}$ where $x_i \in \{0,1\}^n$ for $i \in \{1,2,\cdots,j\}$ and $x_i \neq x_j$
for $i \neq j$. Therefore, the minimum number of the deletion channel outputs is $N_{\ell}(n,t)+1$ in order to reconstruct a list of size $\ell-1$ of candidate sequences, one of which corresponds to the original transmitted sequence. 

Since our results concern the binary case, the first open question is the following.  

\medskip
\noindent
\textbf{Open Question 1.} 
What is the maximum possible size of the intersection of the $t$-deletion balls 
centered at $x_1, x_2, \dots, x_{\ell}$, where 
$x_i \in \{0,1,\dots,q-1\}^n$ for all $i \in \{1,2,\dots,\ell\}$ and 
$x_i \neq x_j$ for $i \neq j$?

\medskip
Since the only condition above is $x_i \neq x_j$ for $i \neq j$, the following 
naturally arises as a second open question.

\medskip
\noindent
\textbf{Open Question 2.} 
What is the maximum possible size of the intersection of the $t$-deletion balls 
centered at $x_1, x_2, \dots, x_{\ell}$, where 
$x_i \in \{0,1,\dots,q-1\}^n$ for all $i \in \{1,2,\dots,\ell\}$ and 
$d_L(x_i, x_j) \geq d$ for $i \neq j$?

\section*{Acknowledgment}
The author would like to thank Prof.~Alexander Barg for suggesting this problem and providing valuable guidance. Additionally, the author extends thanks to  Yihan Zhang for insightful discussions. 

\bibliographystyle{plain}
\bibliography{main.bib}

\end{document}